# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v2

@@ A.1 Proof of Theorem 1
Proof.
Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1).
Using Skorokhod and Yadrenko (1973); Ibragimov and Rozanov (1978); Stein (2004), if for some $a>0$ the function $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some finite and positive $c$,
$\int_{c}^{\infty}z^{d-1}\left(\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right)^{2}\,{\rm d}z<\infty,$ (26)
then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$. Condition (26) is the key tool.
We first need the following result.
@@
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined at (7). Assume $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $2(\beta+\gamma)\geq 6\delta+1$. Then:
@@
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ $=\sigma^{2}Ka^{d}\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
    $\qquad+c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
    as $z\to\infty$;
@@
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta},\qquad z\to\infty.$
Here
$c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},$
$c_{5}=\frac{\pi}{2}\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
We are now ready to prove the theorem.
From Lemma 2 (Point 2), there exist constants $0<c_{1}<c_{2}<\infty$ such that, for $z$ large,
$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)\leq c_{2},\qquad i=0,1.$ (27)
In particular, $\widehat{\rho}_{0}(z)z^{2\delta}$ is bounded away from 0 and $\infty$ as $z\to\infty$.
Using Lemma 2 (Points 1 and 2), and collecting terms, one obtains for large $z$
$\left|\frac{\sigma_{1}^{2}\widehat{\varphi}_{1}(z)-\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}{\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}\right|\leq\frac{1}{c_{1}}\left|A(z)+B(z)+D(z)\right|,$
where
$A(z)$ $=\big[c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big]\{1+\mathcal{O}(z^{-2})\},$
$B(z)$ $=\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z+c_{5}^{1})$
$\quad-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z+c_{5}^{0})\big],$
$D(z)$ $=\mathcal{O}(z^{-1})\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\big].$
Hence condition (26) becomes
$\int_{c}^{\infty}z^{d-1}\big(A(z)+B(z)+D(z)\big)^{2}\,{\rm d}z<\infty.$
Assume now (10), i.e.
$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}}.$
This is equivalent to
$c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}=c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0},$
so that $A(z)\equiv 0$.
Using $(x+y)^{2}\leq 2x^{2}+2y^{2}$, it suffices to check
$\int_{c}^{\infty}z^{d-1+2(\lambda_{i}+2\delta)}\,dz<\infty,\qquad i=0,1,$
which holds iff
@@
$\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta.$
Therefore, under
$\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})>\frac{d+1}{2}+3\delta,$
condition (26) holds and the sufficient part is proved.
The necessary part follows from Zhang (2004).
@@ A.2 Proof of Theorem 2
Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$, $l\geq 0$.
@@
Reduction to $l\geq 0$. The model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$. The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality.
@@ A.3 Proof of Theorem 3
Without loss of generality set $a=1$; the general case follows by the substitution $x\mapsto ax$, which contributes the factor $a^{d}$. Using the change of variable $t=1-x^{2}$ (so $x^{d-1}\,dx=\tfrac{1}{2}(1-t)^{(d-2)/2}\,dt$) in the definition of the integral range (12), and substituting (8), one obtains
$A_{\kappa,\mu,l,1,d}$ $=\frac{\pi^{d/2}C_{0}}{2^{d}\Gamma(d/2)}\int_{0}^{1}(1-t)^{\frac{d-2}{2}}\,t^{\mu+\frac{d-1}{2}+2\kappa}$ (28)
$\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;t\right)\,dt.$
where $C_{0}=\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)\big/\!\left[\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\tfrac{1}{2}\right)\right]$.
We apply formula 2.21.1.11 of Prudnikov et al. (1990),
$\int_{0}^{1}t^{\sigma-1}(1-t)^{\rho-1}\,{}_{2}F_{1}(a,b;c;t)\,dt=\frac{\Gamma(\rho)\Gamma(\sigma)}{\Gamma(\rho+\sigma)}\,{}_{3}F_{2}(a,b,\sigma;\,c,\,\rho+\sigma;\,1),$ (29)
with the identifications
$\sigma=\mu+\tfrac{d+1}{2}+2\kappa,\quad\rho=\tfrac{d}{2},\quad a=\tfrac{\mu}{2},\quad b=\tfrac{\mu+d}{2}+\kappa,\quad c=\mu+\tfrac{d+1}{2}+2\kappa.$
Since $c=\sigma$, the third upper parameter and the first lower parameter of $_{3}F_{2}$ coincide, so the series telescopes term by term:
$_{3}F_{2}(a,b,\sigma;\,\sigma,\,\rho+\sigma;\,1)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}(\sigma)_{k}}{k!\,(\sigma)_{k}(\rho+\sigma)_{k}}=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{k!\,(\rho+\sigma)_{k}}={}_{2}F_{1}(a,b;\,\rho+\sigma;\,1).$
Gauss’s summation theorem then gives, provided $\rho+\sigma-a-b>0$,
$_{2}F_{1}(a,b;\rho+\sigma;1)=\frac{\Gamma(\rho+\sigma)\,\Gamma(\rho+\sigma-a-b)}{\Gamma(\rho+\sigma-a)\,\Gamma(\rho+\sigma-b)}.$
One verifies directly that
$\rho+\sigma-a-b=\kappa+\tfrac{1}{2}+l>0\quad(\kappa>-\tfrac{1}{2},\;l\geq 0),$
$\rho+\sigma-a=\tfrac{\mu+1}{2}+\kappa+d,\qquad\rho+\sigma-b=\tfrac{\mu+1}{2}+\kappa+\tfrac{d}{2}.$
Substituting back into (28), the factors $\Gamma(\sigma)$ and $\Gamma(\rho+\sigma)$ cancel with those arising from (29) and from $C_{0}$. Using the duplication formula $\Gamma(z)\Gamma(z+\tfrac{1}{2})=\sqrt{\pi}\,2^{1-2z}\Gamma(2z)$ to simplify the remaining Gamma factors, one obtains
$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\,\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\,\Gamma\!\left(\kappa+\frac{1}{2}\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa+l\right)},$
which, after restoring the factor $a^{d}$, gives (14). ∎
@@ A.4 Proof of Theorem 4
Without loss of generality, set the compact support to $a=1$. Define
$LA_{\kappa,d}(\mu,l):=\log(A_{\kappa,\mu,l,d,1}),$
@@
$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
We maximize $LA_{\kappa,d}(\mu,l)$ over the admissible set in $\mu$ for a fixed $l\geq 0$, as given by Theorem 2. Let $\delta=\frac{d}{2}+\kappa$. Then
  1. 1.
    $\mu\geq\delta-l+\tfrac{1}{2}$ if $0\leq l\leq\delta$;
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ if $l>\delta$.
This is a maximization problem over the half-line $\mu_{\min}(l),\infty)$.
Let $\psi(x)=\frac{d}{dx}\log\Gamma(x)$ be the digamma function, which is strictly increasing for $x>0$. Set $x=\frac{\mu+1}{2}+\kappa$. Under the validity conditions we have $x>0$ and $x+l>0$. Differentiating with respect to $\mu$ yields
$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi(x+\tfrac{d}{2})+\psi(x+l)-\psi(x+l+\tfrac{d}{2})\Big].$
Since $\psi$ is increasing and $\tfrac{d}{2}>0$, we have $\psi(x)-\psi(x+\tfrac{d}{2})\leq 0$ and $\psi(x+l)-\psi(x+l+\tfrac{d}{2})\leq 0$, hence $\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0$ for all admissible $\mu$. Therefore $LA_{\kappa,d}(\mu,l)$ is non-increasing in $\mu$, and its maximum is achieved at the smallest admissible value of $\mu$, i.e.
$\mu_{1}(l)=\delta-l+\tfrac{1}{2}\quad(0\leq l\leq\delta),\qquad\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l\quad(l>\delta).$
This proves Theorem 4. ∎

### v3

@@ A.1 Proof of Theorem 1
A preliminary Lemma is needed for the proof of Theorem 1. The following result can be obtained from the asymptotic expansion of the generalized hypergeometric function $_{1}F_{2}$ for large argument, as given in Luke (1969).
@@
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined as in (7). Assume $\delta>\frac{d}{2}$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$, and $2(\beta+\gamma)\geq 6\delta+1$. Then:
@@
    As $z\to\infty$,
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=K_{\delta,d}\Big[c_{3}\,(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}+c_{4}\,(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
    where
    $c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},\qquad c_{5}=\frac{\pi}{2}\!\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
@@
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta}$ as $z\to\infty$.
Proof of Theorem 1.
Let
$\rho_{i}(\boldsymbol{h})=\sigma_{i}^{2}\,\varphi_{i}(\|\boldsymbol{h}\|),\qquad\varphi_{i}(x)={\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}}(x),\qquad i=0,1,$
and let $f_{i}$ denote the (isotropic) spectral density of $\rho_{i}$ in the sense of (1). Then $f_{i}(z)=\sigma_{i}^{2}\,\widehat{\varphi}_{i}(z)$, where $\widehat{\varphi}_{i}(z)=\widehat{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)$ is given by (7). Since we are considering valid covariance models, $\widehat{\varphi}_{i}(z)\geq 0$ for all $z\geq 0$, and the Gamma factors in (7) are finite (in particular, $\beta_{i}>d/2$, $\gamma_{i}>d/2$ and $\beta_{i}>\delta$, $\gamma_{i}>\delta$).
We use the standard equivalence criterion for stationary Gaussian measures on bounded domains (see, e.g., Ibragimov and Rozanov (1978) and Stein (2004)): if for some $\eta>0$ the function $f_{0}(z)\,z^{\eta}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some $c>0$
$\int_{c}^{\infty}z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}\,dz<\infty,$ (30)
then $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent on the paths over any bounded $D\subset\mathbb{R}^{d}$. (For $d=1,2,3$ this condition is also sharp in the present regularly-varying setting; see the necessity argument below.)
Set, for $i=0,1$,
$\lambda_{i}:=\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}.$
By assumption, for each $i=0,1$ we have $\delta>\frac{d}{2}$ and $2(\beta_{i}-\delta)(\gamma_{i}-\delta)\geq\delta.$ Moreover, the theorem assumes $\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta,$ hence $2(\beta_{i}+\gamma_{i})>(d+1)+6\delta\geq 6\delta+2>6\delta+1,$ so that in particular $2(\beta_{i}+\gamma_{i})\geq 6\delta+1$. Therefore all assumptions of Lemma 1 hold for $\widehat{\varphi}_{i}$. By Lemma 1 (applied to $\widehat{\varphi}_{i}$), as $z\to\infty$,
$\widehat{\varphi}_{i}(z)=K_{\delta,d}\,L(\beta_{i},\gamma_{i})\,a_{i}^{\,d-2\delta}\,z^{-2\delta}\,\{1+\mathcal{O}(z^{-2})\}\;+\;\mathcal{O}\!\big(z^{\lambda_{i}}\big),$ (31)
where $K_{\delta,d}:=\Gamma(\delta)/(2^{2\delta}\pi^{d/2}\Gamma(\delta-d/2))>0$ depends only on $(\delta,d)$ and
$L(\beta_{i},\gamma_{i})=\frac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}.$
Multiplying by $\sigma_{i}^{2}$, we obtain for $f_{i}(z)=\sigma_{i}^{2}\widehat{\varphi}_{i}(z)$:
$f_{i}(z)=A_{i}\,z^{-2\delta}\{1+\mathcal{O}(z^{-2})\}+\mathcal{O}\!\big(z^{\lambda_{i}}\big),\qquad A_{i}:=\sigma_{i}^{2}\,K_{\delta,d}\,L(\beta_{i},\gamma_{i})\,a_{i}^{\,d-2\delta}.$ (32)
@@
Since the theorem assumes $\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta,$ we have
$\lambda_{i}+2\delta=\frac{1}{2}+3\delta-(\beta_{i}+\gamma_{i})<-\frac{d}{2},$
hence $z^{\lambda_{i}}=o(z^{-2\delta})$ and (32) yields
$f_{i}(z)\sim A_{i}\,z^{-2\delta}\qquad(z\to\infty),\quad\text{with }A_{i}>0.$ (33)
In particular, $f_{0}(z)\,z^{2\delta}$ is bounded away from 0 and $\infty$ for large $z$, so (30) can be applied with $\eta=2\delta$.
Sufficiency. Assume (10). Since $K_{\delta,d}>0$ is common to both models, (10) is equivalent to $A_{0}=A_{1}$ (because $A_{i}$ differs from $\sigma_{i}^{2}L(\beta_{i},\gamma_{i})/a_{i}^{2\delta-d}$ only by the common factor $K_{\delta,d}$). Hence, by (32),
$f_{1}(z)-f_{0}(z)=A_{0}z^{-2\delta}\mathcal{O}(z^{-2})\;+\;\mathcal{O}\!\big(z^{\lambda_{0}}\big)\;+\;\mathcal{O}\!\big(z^{\lambda_{1}}\big).$
Dividing by $f_{0}(z)\sim A_{0}z^{-2\delta}$ gives
$\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}=\mathcal{O}(z^{-2})+\mathcal{O}\!\big(z^{\lambda_{0}+2\delta}\big)+\mathcal{O}\!\big(z^{\lambda_{1}+2\delta}\big).$
Therefore, for large $z$,
$z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}\;\lesssim\;z^{d-1}z^{-4}\;+\;z^{d-1}z^{2(\lambda_{0}+2\delta)}\;+\;z^{d-1}z^{2(\lambda_{1}+2\delta)}.$
The first term is integrable at $\infty$ iff $d<4$, i.e. $d=1,2,3$. For the other terms, note that
$d-1+2(\lambda_{i}+2\delta)=d-1+2\Big(\tfrac{1}{2}+3\delta-(\beta_{i}+\gamma_{i})\Big)=d+6\delta-2(\beta_{i}+\gamma_{i}),$
which is <1<-1 exactly when $\beta_{i}+\gamma_{i}>(d+1)/2+3\delta$, assumed in the theorem. Thus the integral (30) is finite, and by the criterion $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent on bounded $D$.
Necessity: Assume the theorem’s condition $\beta_{i}+\gamma_{i}>(d+1)/2+3\delta$ holds for $i=0,1$ and suppose (10) fails. Then $A_{1}\neq A_{0}$. Using (33),
$\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}=\frac{A_{1}z^{-2\delta}-A_{0}z^{-2\delta}+o(z^{-2\delta})}{A_{0}z^{-2\delta}\{1+o(1)\}}\longrightarrow\frac{A_{1}-A_{0}}{A_{0}}\neq 0\qquad(z\to\infty).$
Hence there exist $c>0$ and $\varepsilon>0$ such that for all $z\geq c$,
$\left|\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right|\geq\varepsilon,$
and therefore
$\int_{c}^{\infty}z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}dz\;\geq\;\varepsilon^{2}\int_{c}^{\infty}z^{d-1}\,dz=\infty.$
Thus (30) fails and the measures cannot be equivalent. Consequently, equivalence implies $A_{1}=A_{0}$, which is exactly (10).
Combining the sufficiency and necessity arguments proves the claimed “if and only if” statement. ∎
@@ A.2 Proof of Theorem 2
Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$.
Note that the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$. The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality. We henceforth assume $l\geq 0$.
@@ A.3 Proof of Theorem 3
Since $\phi(\|\cdot\|)$ is compactly supported it is absolutely integrable and, by the spectral representation associated with (1),
$A=(2\pi)^{d}\,\widehat{\phi}(0).$
For $\phi(x)=\mathcal{GH}_{\delta,\beta,\gamma,a}(x)$, the spectral density is (7). Evaluating at $z=0$ and using $_{1}F_{2}(\cdot;\,0)=1$ gives
$\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(0)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}.$
Hence
$A=(2\pi)^{d}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(0)=\pi^{d/2}a^{d}\,\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}.$
Substituting $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$ yields (14). ∎
@@ A.4 Proof of Theorem 4
Since $A_{\kappa,\mu,l,a,d}=a^{d}\,A_{\kappa,\mu,l,1,d}$, the factor $a^{d}$ is irrelevant for monotonicity and maximizers in $\mu$. Hence we set $a=1$.
Fix $l\geq 0$ and consider
$LA_{\kappa,d}(\mu,l):=\log\!\big(A_{\kappa,\mu,l,1,d}\big),$
@@
$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{\Gamma\!\left(\kappa+\frac{1}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
Let $\psi=\Gamma^{\prime}/\Gamma$ be the digamma function and set
$x=\frac{\mu+1}{2}+\kappa\qquad\Big(\text{so }\frac{dx}{d\mu}=\frac{1}{2}\Big).$
On the validity region of the model (hence in particular on the admissible domain in $\mu$), the Gamma factors above are well-defined and positive, so $x>0$ and $x+l>0$. Differentiating with respect to $\mu$ gives
$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi\!\left(x+\tfrac{d}{2}\right)+\psi(x+l)-\psi\!\left(x+l+\tfrac{d}{2}\right)\Big].$
Since $\psi$ is strictly increasing on $(0,\infty)$ and $d/2>0$, each difference $\psi(x)-\psi(x+d/2)$ and $\psi(x+l)-\psi(x+l+d/2)$ is $\leq 0$. Therefore
$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0\quad\text{for all admissible }\mu,$
hence $\mu\mapsto A_{\kappa,\mu,l,a,d}$ is non-increasing on its admissible domain. Consequently, its maximum is attained at the smallest admissible value of $\mu$.
We now identify this smallest admissible value using Theorem 2. If $0\leq l\leq d/2+\kappa$, Theorem 2 provides the exact validity condition $\mu\geq\mu_{1}(l):=\delta-l+\frac{1}{2}$, hence $\mu^{\star}(l)=\mu_{1}(l)$.
If $l>d/2+\kappa$, Theorem 2 guarantees validity for all $\mu\geq\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l$. In this regime, we take as admissible domain the parameter range ensured by Theorem 2 (i.e., $\mu\geq\mu_{2}(l)$), and by monotonicity the maximum over this admissible domain is attained at $\mu^{\star}(l)=\mu_{2}(l)$. ∎